OpenStudy (anonymous):

Simplify the complex fraction: @kx2bay

OpenStudy (anonymous):

\[\frac{ \frac{ \frac{ 1 }{ a^2+8a+12 } }{ 1 } }{ a+2 }\]

OpenStudy (anonymous):

\[\frac{ \frac{ \frac{ 1 }{ (a+2)(a+6) } }{ 1 } }{ a+2 }\]

OpenStudy (anonymous):

the a+2s cancel out

OpenStudy (anonymous):

correct cookie

OpenStudy (anonymous):

and u left with 1 / (a+6) , right? where u from btw?

OpenStudy (anonymous):

America, the state of California :)

OpenStudy (anonymous):

oh cool, lots of oranges there ;)

OpenStudy (anonymous):

lol u mean the fruit or something else lol cuz white people tan & they look orange lol

OpenStudy (anonymous):

hahahahaha , no i meant the fruit, we get it here in oz land from california

OpenStudy (anonymous):

really? we're popular because of oranges? lol never knew

OpenStudy (anonymous):

and the beaches of course...

OpenStudy (anonymous):

Oh, we have so many beaches lol Don't forget Disneyland lol We have 2 parks in one, kinda lol

OpenStudy (anonymous):

hollywood too, u guys very lucky, got all the action..

OpenStudy (anonymous):

hollywood too, u guys very lucky, got all the action..

OpenStudy (anonymous):

We call it HollyWeird...weird pellet & people be in Hollywood lol I barely go myself

OpenStudy (anonymous):

:)